Sales revenue Cost of goods sold (all variable) Gross margin Selling expenses (20% variable) Administrative expenses (60% variable) Operating income $ 56,632 28,967 27,665 8,905 12,800 $ 5,960 v (a Prepare a contribution format income statement for Mary, (Round answers to o decimal places, e.g. 5,250.)
Answer #1
Dear Student
Thank you for using homeworklib
Please find below the answer
Statementshowing Computations
Paticulars Amount Amount
Sales revenues              56,632.00
Variable costs:
Variable cost of goods sold =28,967*100%        28,967.00
Variable selling expenses =8905*20%           1,781.00
Variable admin expenses = 12800*60%           7,680.00
Total variable costs              38,428.00
Contribution              18,204.00
Fixed cost of goods sold                        -  
Fixed selling expenses=8905*80%           7,124.00
Fixed admin expenses =12800*40%           5,120.00
Total Fixed costs              12,244.00
Operating income                 5,960.00